how much money must you deposit into a savings account that has a 3% annual interest rate to earn 100.00 in three years?
Answer by mananth(16946)

how much money must you deposit into a savings account that has a 3% annual interest rate to earn 100.00 in three years?
...
Simple Interest = p*r*t
100=p*0.03*3
p=100/3*0.03
p=100/0.09
p= $1111.11